1.1 The speed of a shaft increases from 93 r/min to 305 r/min with an angular acceleration of 2,4 rad/s2. How many revolutions does the shaft make during the first 3 seconds of acceleration? How many revolutions does the shaft make in the last 4.25seconds. Show all calculations[10]

1.2 A motor that runs at 1240 rev/minute is required to drive a belt at 20 m/s. What diameter belt pulley is required?[3]


Also clearly show the following values:

ωi

ωf

θ

r

d

ωf

1
Expert's answer
2021-08-19T02:21:03-0400

1.1 Let us find out how many revolutions does the shaft make during the first 3 seconds of acceleration:


"n(t)=n_it+\frac{\alpha t^2}{2},\\\space\\
n_f=n(3)=(93/60)3+\frac{(2.4/(2\pi)) ·3^2}{2}=6."

Now, find the total time required to increase the speed from 93 to 305 r/min:


"n(T)=n_0T+\frac{\alpha T^2}{2},\\\space\\
n(T)=(305/60)\text{ r/s},\\\space\\
(305/60)=(93/60)T+\frac{(2.4/(2\pi))·T^2}{2},\\\space\\
0.191T^2+1.55T-5.083=0,\\\space\\
T=\frac{-1.55+\sqrt{1.55^2-4·0.191·(-5.083)}}{2·0.191}=5.013\text{ s}."

Therefore, the time before the last 4.25 s is


"\tau=5.013-4.25=0.76\text{ s}."

How many revolutions during the 0.76 s:


"n(0.76)=(93/60)0.76+\frac{(2.4/(2\pi)) ·0.76^2}{2}=1."

How many revolutions during the 5.013 s:


"n(5.013)=(93/60)5.013+\frac{(2.4/(2\pi)) ·5.013^2}{2}=12.\\\space\\
n_\text{last 4.25 s}=n(5.013)-n(0.76)=11."

1.2 Find the angular velocity:


"\omega=2\pi n."

The velocity:


"v=\omega R=2\pi n R,\\\space\\
R=\frac{v}{2\pi n}=\frac{20}{2\pi·(1240/60)}=0.15\text{ m}."